OpenStudy (anonymous):

Nora's savings account has a balance of $3979. After 4 years, what will the amount of interest be at 12% compounded semiannually?

OpenStudy (anonymous):

2(3979*0.12)+3979 does this help

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2(3979*0.12)+3979 2(477.48)+3979 954.96+3979 4933.96

OpenStudy (anonymous):

still didnt help its not one of my answers

OpenStudy (mathstudent55):

\(\large F = P(1 + \dfrac{r}{n})^{nt}\) F = future value P = present value r = annual interest rate as decimal n = number of compounding periods in one year t = number of years

OpenStudy (mathstudent55):

\(F = $3979(1 + \dfrac{0.12}{2})^{(2)4} \)

OpenStudy (mathstudent55):

What are your choices?

OpenStudy (anonymous):

$2362.92 this wat i got

OpenStudy (anonymous):

a. $2367.92 c. $2362.92 b. $4774.80 d. $2353.92

OpenStudy (mathstudent55):

How can you end up with less than you started with?